OpenStudy (anonymous):

The base of a pyramid is a square with dimensions 12 in by 12 in. The height of the pyramid is 10 in. What is the slant of the pyramid

OpenStudy (anonymous):

Damn, I can't remember the formula for this. Let me look it up...

OpenStudy (anonymous):

s² = h² + (b/2)²

OpenStudy (anonymous):

s^2 = 10^2 + (144/2)^2

OpenStudy (anonymous):

s^2 = 100 + 72^2

OpenStudy (jtvatsim):

correction s^2 = 10^2 + (12/2)^2 b = 12 not 144. :)

OpenStudy (anonymous):

derp

OpenStudy (jtvatsim):



OpenStudy (anonymous):

jvat, are you sure? I thought that the base of a pyramid was l*w

OpenStudy (jtvatsim):

well, don't get caught up on the formula and focus on the geometry of the question

OpenStudy (imstuck):

The picture is correct. Now to find the slant height you use Pythagorean's Theorem and solve for the hypotenuse.

OpenStudy (jtvatsim):

the question has to do with lengths, so introducing area is not appropriate, the units would be wrong.

OpenStudy (anonymous):

Wouldn't it come out the same way?

OpenStudy (jtvatsim):

let's try it and see. :)

OpenStudy (imstuck):

Just draw the pic using the info given and it is apparent what you need to do to solve this.

OpenStudy (imstuck):

What did both of you get as the measure of the slant height?

OpenStudy (jtvatsim):

11.66

OpenStudy (imstuck):

And where is Edelong in all of this discussion, I wonder? lol

OpenStudy (jtvatsim):

ikr

OpenStudy (imstuck):

I got that too. (11.66)

OpenStudy (anonymous):

My result is different. I didn't see it the way you saw it. :/

OpenStudy (jtvatsim):

the problem here was in interpreting the formula

OpenStudy (jtvatsim):

which is why I try to not trust formulas, it is important to understand what the situation is actually about. :)